# DAS_CE_Supp

### v2

@@ IV-A Observation Matrix Design Using Greedy Method
Figure 2: Framework of the proposed channel estimation method.
In practical uplink channel estimation where the noise vector ${\mathbf{z}}$ is amplified by the observation vector ${\mathbf{w}}_{q}$, the pilot-wise power constraint should be considered. In this case, the MIM problem in (6) is an intractable non-convex programming. To overcome this challenge, we harness a greedy method to obtain the observation vectors in a pilot-by-pilot manner, as shown in Fig. 2. Define ${\bf{W}}_{t}=[{\bf{w}}_{1},{\bf{w}}_{2},\cdots,{\bf{w}}_{t}]$ as the observation matrix for timeslots $1\sim t$, where $t<Q$, and denote ${\mathbf{y}}_{t}={\mathbf{W}}_{t}^{H}{\mathbf{h}}+{\mathbf{z}}_{t}$ as the corresponding received signal. Given the current observation matrix ${\bf{W}}_{t}$, our greedy method aims to determine the next observation vector, ${\mathbf{w}}_{t+1}$, by maximizing the MI increment from timeslot $t$ to $t+1$, i.e., $\max_{{\mathbf{w}}_{t+1}}I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({\mathbf{y}}_{t};{\mathbf{h}})$. As proved in [49], the MI increment is expressed as
@@
where ${{\bf{\Sigma}}_{t}}$ represents the posterior kernel matrix of ${\mathbf{h}}$ given the observation ${\mathbf{y}}_{t}$, i.e.,
Thereafter, the optimal ${\bf{w}}_{t+1}$ can be attained by solving the quadratic problem:
@@
Before delving into the detail this problem, it is notable that repeatedly calculating the inverse of ${{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ in each timeslot $t$ is necessary for obtaining the kernel ${\bf{\Sigma}}_{t}$ in (9) and (10). To avoid these computationally inefficient matrix-inversion operations, Lemma 1 would be helpful to offer an efficient approach to obtain ${\bf{\Sigma}}_{t}$.
Lemma 1
${\bf{\Sigma}}_{t}$ can be calculated from ${\bf{\Sigma}}_{t-1}$ and ${\mathbf{w}}_{t}$ as:
Proof:
See Appendix A. ∎
Clearly, Lemma 1 reveals the relationship between ${\bf{\Sigma}}_{t}$ and ${\bf{\Sigma}}_{t-1}$, and thereby allowing us to update ${\bf{\Sigma}}_{t}$ from ${\bf{\Sigma}}_{t-1}$ without the need for matrix inversion.
As a result, equations (8) to (11) inspire us to propose the framework presented in Fig. 2 for the design of observation matrix. The framework begins with setting ${\bm{\Sigma}}_{0}$ as the prior kernel of the channel, ${\bm{\Sigma}}_{\bf h}$. The first observation vector ${\mathbf{w}}_{1}$ is obtained by solving the quadratic problem in (10). For subsequent timeslots $t$ ($t\in\{1,2,\cdots,Q-1\}$), the posterior kernels ${\bm{\Sigma}}_{t}$ are updated from ${\bm{\Sigma}}_{t-1}$ and ${\bf w}_{t}$ using Lemma 1. The update of the posterior kernels then triggers the update of the observation vectors ${\bf w}_{t+1}$ by solving problem (10) pilot-by-pilot. These two updating steps are repeated until all $Q$ observation vectors are generated for performing the ensuing Bayesian regression.
@@ IV-B Ice-Filling for Amplitude-and-Phase Controllable Observation Matrix
2:  Initialize: $[\lambda_{1}^{0},\lambda_{2}^{0},\cdots,\lambda_{K}^{0}]=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$
@@
7:     Eigenvalue-preserve: $\lambda_{k}^{t+1}=\lambda_{k}^{t}$ with $k\in\{1,2,\cdots,K\}/\{k_{t}\}$
@@
where the principal eigenvector of ${\bf{\Sigma}}_{t}$ yields the optimal solution to ${\mathbf{w}}_{t+1}$, i.e., ${\bf{\Sigma}}_{t}{\bf{w}}_{t+1}=\lambda_{\max}({\bf{\Sigma}}_{t}){\bf{w}}_{t+1}$. By leveraging the updating strategy of ${\bm{\Sigma}}_{t}$ described in Lemma 1 and the property of eigenvector, we can prove Theorem 1, which lays the foundation for our ice-filling algorithm.
@@
Applying the property of the principal eigenvector ${\bf\Sigma}_{t-1}{\mathbf{w}}_{t}=\lambda_{\max}({\bm{\Sigma}}_{t-1}){\mathbf{w}}_{t}$, we get ${\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}=\lambda_{\max}({\bf\Sigma}_{t-1})$ and ${\bf{\Sigma}}_{t-1}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}=\lambda_{\max}^{2}({\bf\Sigma}_{t-1}){\mathbf{w}}{\mathbf{w}}^{H}$, which together with Lemma 1 give rise to (13). ∎
@@
Given that ${\mathbf{w}}_{t}$ is the principal eigenvector of ${\bm{\Sigma}}_{t-1}$, Theorem 1 implies that the posterior kernels, ${\bm{\Sigma}}_{t}$ and ${\bm{\Sigma}}_{t-1}$, share the identical eigenspace. This fact immediately leads to the result that, for the amplitude-and-phase controllable observation matrix, the eigenvectors, $\{{\mathbf{u}}_{k}\}$, of the prior kernel, ${\bm{\Sigma}}_{0}={\bm{\Sigma}}_{{\mathbf{h}}}$, are inherited by all subsequent posterior kernels, ${\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2},\cdots,$ and ${\bm{\Sigma}}_{Q-1}$. Thereby, we can draw the conclusion that all observation vectors, $\{{\mathbf{w}}_{t}\}$, are picked from the eigenvectors of ${\bm{\Sigma}}_{{\mathbf{h}}}$.
@@
In terms of the eigenvalues when updating ${\bm{\Sigma}}_{t-1}$ to ${\bm{\Sigma}}_{t}$, it can be straightforwardly proved from Theorem 1 that only the principal eigenvalue, $\lambda_{\max}({\bm{\Sigma}}_{t-1})$, of ${\bm{\Sigma}}_{t-1}$ is squeezed to an eigenvalue of ${\bm{\Sigma}}_{t}$ given by
$\lambda_{\max}({\bm{\Sigma}}_{t-1})-\frac{\lambda_{\max}^{2}({\bf\Sigma}_{t-1})}{\lambda_{\max}({\bf\Sigma}_{t-1})+\sigma^{2}}=\frac{\lambda_{\max}({\bf\Sigma}_{t-1})\sigma^{2}}{\lambda_{\max}({\bf\Sigma}_{t-1})+\sigma^{2}},$ (14)
while the other eigenvalues are preserved.
As a result, updating kernels in amplitude-and-phase controllable situations is equivalent to squeezing the eigenvalues within the identical eigenspace.
Motivated by Theorem 1, our proposed ice-filling algorithm is intrinsically an eigenvector-assignment-process, as presented in Algorithm 1. To elaborate, we define $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ as the eigenvalues of the posterior kernel ${\bm{\Sigma}}_{t}$, which are initialized as the eigenvalues of the prior kernel ${\bm{\Sigma}}_{\bf h}$ in Step 2. In each timeslot, we find the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ and index it by $k_{t}$ in Step 4. Then, the corresponding eigenvector ${\mathbf{u}}_{k_{t}}$ of ${\bm{\Sigma}_{h}}$, which is equivalent to the principal eigenvector of ${\bm{\Sigma}}_{t}$, is assigned to ${\mathbf{w}}_{t+1}$ in Step 5. Finally, according to Theorem 1, the selected eigenvalue $\lambda_{k_{t}}^{t}$ is squeezed to $\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}$, while the other eigenvalues of ${\bm{\Sigma}}_{t}$ are preserved, to attain the eigenvalues of the next kernel ${\bm{\Sigma}}_{t+1}$. These steps are repeatedly executed until all observation vectors are generated, which completes the ice-filling algorithm.
@@ IV-C Ice-Filling Versus Water-Filling
In this subsection, we put forward a more insightful interpretation to the ice-filling algorithm that establishes its connection with the water-filling principle. Generally speaking, the proposed ice-filling algorithm can be viewed as a quantization of the water-filling principle. Specifically, it is clear that the observation matrices generated by both the water-filling and the ice-filling algorithms fall within the eigenspace of ${\bm{\Sigma}}_{\bf h}$. The major distinction is attributed to their “power allocation strategies”. The water-filling principle allocates $Q$ units of power to the $K$ eigenvectors using (7). On the other hand, the ice-filling algorithm transforms this power-allocation process into an eigenvector-assignment process, where the $K$ eigenvectors are repeatably assigned to the observation vectors in $Q$ timeslots.
@@
Comparing the definitions of $p_{k}$ and $n_{k}$, we can intuitively interpret the $n_{k}$-timeslot reuse of the observation vector ${\mathbf{u}}_{k}$ for pilot transmission as allocating $n_{k}$ units of power to ${\mathbf{u}}_{k}$. More importantly, Definition 1 enables us to derive the analytical expressions of the eigenvalues $\{\lambda_{k}^{t}\}$ in Algorithm 1. Note that the eigenvalue update rule of the ice-filling algorithm can be rewritten as a recursive formula:
@@
Interpretation of ice-filling: As depicted in Fig. 3, the water-filling principle allocates $Q$ units of water (power) to a vessel with $K$ channels, each having a unique base level $\frac{\sigma^{2}}{\lambda_{k}}$, $k\in\{1,2,\cdots,K\}$. By controlling the uniform water level $\beta$, the optimal power $p_{k}$ is determined by the gap between the base level and the water level. In contrast, the ice-filling algorithm transforms the total $Q$ units of water into $Q$ ice blocks, each containing one unit of power and being used for one pilot transmission. Our algorithm starts from an empty vessel and fills one ice block onto the channel having the deepest base surface $\frac{\sigma^{2}}{\lambda_{k_{0}}}$. This operation is equivalent to finding the largest eigenvalue $\lambda_{k_{0}}$ from $\{\lambda_{1},\cdots,\lambda_{K}\}$. Then, the eigenvector ${\mathbf{u}}_{k_{0}}$ is assigned to the first observation vector ${\mathbf{w}}_{1}$, and ice level of this channel increases from $\frac{\sigma^{2}}{\lambda_{k_{0}}}$ to $\frac{\sigma^{2}}{\lambda_{k_{0}}^{1}}=\frac{\sigma^{2}}{\lambda_{k_{0}}}+1$. In the subsequent time slots, the remaining $Q-1$ ice blocks are filled onto the channels locating at the deepest base surfaces or ice levels, indexed by $k_{t}=\arg\min_{k}\{\frac{\sigma^{2}}{\lambda_{k}^{t}}\}$, one by one. The corresponding eigenvectors ${\mathbf{u}}_{k_{t}}$ are used for pilot transmission, and the ice levels increase according to (16). Consequently, the final pilot reuse frequencies $\{n_{k}\}$ are determined by the number of ice blocks on top of each channel, and the ice levels are given by $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}=\{n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\}$.
The comparison between the ice-filling and water-filling in Fig. 3 indicates that the continuous powers $\{p_{k}\}$ are quantized into the discrete pilot reuse frequencies $\{n_{k}\}$. This quantization results in the non-uniform ice levels $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}$, as opposed to the uniform water level $\beta$. To see the quantization nature of ice-filling more clearly, we prove Theorem 2 to upper bound the quantization error between $p_{k}$ and $n_{k}$.

### v3

@@ IV-A Observation Matrix Design Using Greedy Method
Figure 2: Framework of the proposed observation matrix design.
In practical uplink channel estimation where the noise vector ${\mathbf{z}}$ is amplified by the observation vector ${\mathbf{w}}_{q}$, the pilot-wise power constraint should be considered. In this case, it is intractable to find the optimal solution to (6). To overcome this challenge, we harness a greedy method to obtain the observation vectors in a pilot-by-pilot manner, as shown in Fig. 2. Define ${\bf{W}}_{t}=[{\bf{w}}_{1},{\bf{w}}_{2},\cdots,{\bf{w}}_{t}]$ as the observation matrix for timeslots $1\sim t$, where $t<Q$, and denote ${\mathbf{y}}_{t}={\mathbf{W}}_{t}^{H}{\mathbf{h}}+{\mathbf{z}}_{t}$ as the corresponding received signal. Given the current observation matrix ${\bf{W}}_{t}$, our greedy method aims to determine the next observation vector, ${\mathbf{w}}_{t+1}$, by maximizing the MI increment from timeslot $t$ to $t+1$, i.e., $\max_{{\mathbf{w}}_{t+1}}I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({\mathbf{y}}_{t};{\mathbf{h}})$. As proved in [49], the MI increment is expressed as
@@
where ${{\bf{\Sigma}}_{t}}$ represents the posterior kernel of ${\mathbf{h}}$ given the observation ${\mathbf{y}}_{t}$. Using the block matrix inverse, we can explicitly write ${\bf{\Sigma}}_{t}$ as a function of ${\bf{\Sigma}}_{t-1}$ and ${\mathbf{w}}_{t}$ (Detailed proof can be found in Appendix A):
Accordingly, the optimal ${\bf{w}}_{t+1}$ can be attained by solving the quadratic problem:
@@
As a result, equations (8) to (10) inspire us to propose the framework in Fig. 2 for designing observation matrices. It begins by setting ${\bm{\Sigma}}_{0}$ as the prior kernel, ${\bm{\Sigma}}_{\bf h}$. Then, the first observation vector ${\mathbf{w}}_{1}$ is obtained by solving problem (10). In subsequent timeslots $t\in\{1,2,\cdots,Q-1\}$, the posterior kernels ${\bm{\Sigma}}_{t}$ are updated from ${\bm{\Sigma}}_{t-1}$ and ${\bf w}_{t}$ using (9), while the observation vectors ${\bf w}_{t+1}$ are updated from ${\bm{\Sigma}}_{t}$ by solving problem (10). These two updating steps are repeated until all $Q$ observation vectors are generated for performing the ensuing Bayesian regression.
@@ IV-B Ice-filling for Amplitude-and-phase Controllable Observation Matrix
2:  Initialize $[\lambda_{1}^{0},\lambda_{2}^{0},\cdots,\lambda_{K}^{0}]=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$
@@
7:     Eigenvalue-preserve: $\lambda_{k}^{t+1}=\lambda_{k}^{t}$ for $k\neq k_{t}$
@@
The principal eigenvector of ${\bf{\Sigma}}_{t}$ yields the optimal solution to ${\mathbf{w}}_{t+1}$, i.e., ${\bf{\Sigma}}_{t}{\bf{w}}_{t+1}=\lambda_{\max}({\bf{\Sigma}}_{t}){\bf{w}}_{t+1}$. By leveraging the updating strategy of ${\bm{\Sigma}}_{t}$ described in (9) and the property of eigenvector, we can prove Theorem 1, which lays the foundation for our ice-filling algorithm.
@@
Applying the property of the principal eigenvector ${\bf\Sigma}_{t-1}{\mathbf{w}}_{t}=\lambda_{\max}({\bm{\Sigma}}_{t-1}){\mathbf{w}}_{t}$, we get ${\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}=\lambda_{\max}({\bf\Sigma}_{t-1})$ and ${\bf{\Sigma}}_{t-1}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}=\lambda_{\max}^{2}({\bf\Sigma}_{t-1}){\mathbf{w}}_{t}{\mathbf{w}}^{H}_{t}$, which together with (9) give rise to (12). ∎
@@
Given that ${\mathbf{w}}_{t}$ is the principal eigenvector of ${\bm{\Sigma}}_{t-1}$, Theorem 1 implies that the posterior kernels, ${\bm{\Sigma}}_{t}$ and ${\bm{\Sigma}}_{t-1}$, share the identical eigenspace. This fact immediately leads to the result that the eigenvectors, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$, of the prior kernel, ${\bm{\Sigma}}_{0}={\bm{\Sigma}}_{{\mathbf{h}}}$, are inherited by all subsequent posterior kernels, ${\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2},\cdots,$ and ${\bm{\Sigma}}_{Q-1}$. Thereby, we can draw the conclusion that all observation vectors, $\{{\mathbf{w}}_{1}\}_{t=1}^{Q}$, are selected from the eigenvectors of ${\bm{\Sigma}}_{{\mathbf{h}}}$.
@@
Consider the eigenvalues of ${\bf\Sigma}_{t-1}$ and ${\bf\Sigma}_{t}$. We denote the EVD of ${\bf\Sigma}_{t-1}$ as ${\bf\Sigma}_{t-1}={\mathbf{U}}_{K}{\rm diag}\{\lambda_{1}^{t-1},\lambda_{2}^{t-1},\cdots,\lambda_{K}^{t-1}\}{\mathbf{U}}_{K}^{H}$, where $\{\lambda_{k}^{t-1}\}_{k=1}^{K}$ refers to the eigenvalues of $\{{\mathbf{u}}_{k}\}_{k=1}^{K}$, respectively. We use $k_{t-1}$ to index the principal eigenvalue of ${\bf\Sigma}_{t-1}$. Accordingly, the equations ${\mathbf{w}}_{t}={\mathbf{u}}_{k_{t-1}}$ and $\lambda_{\max}({\bf\Sigma}_{t-1})=\lambda_{k_{t-1}}^{t-1}$ hold, and Theorem 1 can be rewritten as:
${\bf{\Sigma}}_{t}$ $={\mathbf{U}}_{K}{\rm diag}\{\lambda_{1}^{t-1},\cdots,\lambda_{k_{t-1}}^{t-1},\cdots,\lambda_{K}^{t-1}\}{\mathbf{U}}_{K}^{H}$
$\quad\quad\quad\quad\quad\quad-\frac{(\lambda_{k_{t-1}}^{t-1})^{2}}{\lambda_{k_{t-1}}^{t-1}+\sigma^{2}}{\mathbf{u}}_{k_{t-1}}{\mathbf{u}}_{k_{t-1}}^{H}$ (13)
$={\mathbf{U}}_{K}{\rm diag}\left\{\lambda_{1}^{t-1},\cdots,\frac{\lambda_{k_{t-1}}^{t-1}\sigma^{2}}{\lambda_{k_{t-1}}^{t-1}+\sigma^{2}},\cdots,\lambda_{K}^{t-1}\right\}{\mathbf{U}}_{K}^{H}.$
Equation (2) reveals that only the principal eigenvalue, $\lambda_{k_{t-1}}^{t-1}$, of ${\bm{\Sigma}}_{t-1}$ is squeezed to an eigenvalue of ${\bm{\Sigma}}_{t}$ given by $\frac{\lambda_{k_{t-1}}^{t-1}\sigma^{2}}{\lambda_{k_{t-1}}^{t-1}+\sigma^{2}}$, while the remaining eigenvalues, $\lambda_{k}^{t-1},\forall k\in\{1,2,\cdots,K\}/\{k_{t-1}\}$, keep unchanged.
Motivated by Remark 1 and Remark 2, our ice-filling algorithm is intrinsically an eigenvector-assignment-process, as presented in Algorithm 1. To elaborate, the eigenvalues $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ of the posterior kernel ${\bm{\Sigma}}_{t}$ are initialized as the eigenvalues of the prior kernel ${\bm{\Sigma}}_{\bf h}$ at $t=0$ in Step 2. In each timeslot, we find the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ and index it by $k_{t}$ in Step 4. Then, the corresponding eigenvector ${\mathbf{u}}_{k_{t}}$ of ${\bm{\Sigma}_{h}}$, which is equivalent to the principal eigenvector of ${\bm{\Sigma}}_{t}$, is assigned to ${\mathbf{w}}_{t+1}$ in Step 5. Finally, according to Remark 2, the selected eigenvalue $\lambda_{k_{t}}^{t}$ is squeezed to $\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}$, while the other eigenvalues of ${\bm{\Sigma}}_{t}$ are preserved, to attain the eigenvalues of the next kernel ${\bm{\Sigma}}_{t+1}$. These steps are repeatedly executed until all observation vectors are generated, which completes the ice-filling algorithm.
@@ IV-C Ice-Filling Versus Water-Filling
In this subsection, we put forward a more insightful interpretation to the ice-filling algorithm that establishes its connection with the water-filling principle. Generally speaking, the proposed ice-filling algorithm can be viewed as a quantization of the water-filling principle. Specifically, it is clear that the observation matrices generated by both the water-filling and the ice-filling algorithms fall within the eigenspace of ${\bm{\Sigma}}_{\bf h}$. The major distinction is attributed to their “power allocation strategies”. The water-filling principle allocates $Q$ units of power to the $K$ eigenvectors using (7). The ice-filling algorithm transforms this power-allocation process into an eigenvector-assignment process, where the $K$ eigenvectors are repeatably assigned to the observation vectors in $Q$ timeslots.
@@
Comparing the definitions of $p_{k}$ and $n_{k}$, we can intuitively interpret the $n_{k}$-timeslot reuse of the observation vector ${\mathbf{u}}_{k}$ for pilot transmission as allocating $n_{k}$ units of power to ${\mathbf{u}}_{k}$. More importantly, Definition 1 enables us to derive the analytical expressions of the eigenvalues $\{\lambda_{k}^{t}\}_{k=1}^{K}$ in Algorithm 1. Note that the eigenvalue update rule of the ice-filling algorithm can be rewritten as a recursive formula:
@@
Interpretation of ice-filling: As depicted in Fig. 3, the water-filling principle allocates $Q$ units of water (power) to a vessel with $K$ channels, each having a unique base level $\frac{\sigma^{2}}{\lambda_{k}}$, $k\in\{1,2,\cdots,K\}$. By controlling the uniform water level $\beta$, the optimal power $p_{k}$ is determined by the gap between the base level and the water level. In contrast, the ice-filling algorithm transforms the total $Q$ units of water into $Q$ ice blocks, each containing one unit of power and being used for one pilot transmission. Our algorithm starts from an empty vessel and fills one ice block onto the channel having the deepest base surface $\frac{\sigma^{2}}{\lambda_{k_{0}}}$. This operation is equivalent to finding the largest eigenvalue $\lambda_{k_{0}}$ from $\{\lambda_{1},\cdots,\lambda_{K}\}$. Then, the eigenvector ${\mathbf{u}}_{k_{0}}$ is assigned to the first observation vector ${\mathbf{w}}_{1}$, and ice level of this channel increases from $\frac{\sigma^{2}}{\lambda_{k_{0}}}$ to $\frac{\sigma^{2}}{\lambda_{k_{0}}^{1}}=\frac{\sigma^{2}}{\lambda_{k_{0}}}+1$. In the subsequent time slots, the remaining $Q-1$ ice blocks are filled onto the channels locating at the deepest base surfaces or ice levels, indexed by $k_{t}=\arg\min_{k}\{\frac{\sigma^{2}}{\lambda_{k}^{t}}\}_{k=1}^{K}$, one by one. The corresponding eigenvectors ${\mathbf{u}}_{k_{t}}$ are used for pilot transmission, and the ice levels increase according to (15). Consequently, the final pilot reuse frequencies $\{n_{k}\}_{k=1}^{K}$ are determined by the number of ice blocks on top of each channel, and the ice levels are given by $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}_{k=1}^{K}=\{n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\}_{k=1}^{K}$.
The comparison between the ice-filling and water-filling in Fig. 3 indicates that the continuous powers $\{p_{k}\}_{k=1}^{K}$ are quantized into the discrete pilot reuse frequencies $\{n_{k}\}_{k=1}^{K}$. This quantization results in the non-uniform ice levels $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}_{k=1}^{K}$, as opposed to the uniform water level $\beta$. To see the quantization nature of ice-filling more clearly, we prove Theorem 2 to upper bound the quantization error between $p_{k}$ and $n_{k}$.